Find f(x - 2) when f(x) =
2x - 5
over
3x + 4

Found 2 solutions by stanbon, unlockmath:
Answer by stanbon(75887)

Find f(x - 2) when f(x) =
2x - 5
over
3x + 4
==============
f(x) = (2x-5)/(3x+4)
----
f(x-2) = (2(x-2)-5)/(3(x-2)+4)
---
= (2x-4-5)/(3x-6+4)
---
= (2x-9)/(3x-2)
===================
Cheers,
Stan H.
===================

Answer by unlockmath(1688)

Hello,
Let's write it as this: f(x - 2) when f(x)=(2x - 5)/(3x+4)
Now substitute (x-2) for each x:
f(x) =(2(x-2)-5)/(3(x-2) + 4)
Expand this and combine like terms to get:
f(x)=(2x-9)/(3x-2)
Make sense?
RJ
www.math-unlock.com